# Learning Quadruped Locomotion Using Differentiable Simulation

### v2

@@ III-D Backpropagation Through Time
$\frac{\partial\mathcal{L}_{\theta}}{\partial\theta}$ $=\frac{1}{N}\sum_{k=0}^{N-1}\left(\sum_{i=1}^{k}\frac{\partial l_{k}}{\partial x_{k}}\prod_{j=i}^{k}\left(\frac{\partial x_{j}}{\partial x_{j-1}}\right)\frac{\partial x_{i}}{\partial\theta}+\frac{\partial l_{k}}{\partial u_{k}}\frac{\partial u_{k}}{\partial\theta}\right),$ (4)
@@
where the matrix of partial derivatives $\partial x_{j}/\partial x_{j-1}$ is the Jacobian of the dynamical system $f$. Therefore, we can compute the policy gradient directly by backpropagating through the differentiable physics model and a loss function $l_{k}$ that is differentiable with respect to the system state and control inputs. A graphical model for gradient backpropagation in policy learning using differentiable simulation is given in Fig. 1. Due to the usage of multiplication $\prod$, there are two potential issues in using Eq. (4) for policy gradient: 1) gradient vanishing or exploding, 2) long computation time. We tackle these two problems via short-horizon policy training.
III-E Short-Horizon Policy Training
Although smoothed physical models (e.g., single rigid-body dynamics) improve the local optimization landscape, the complexity of the optimization problem escalates significantly in long-horizon problems involving extensive concatenation of simulation steps. The situation further deteriorates when the actions within each step are interconnected through a nonlinear and nonconvex neural network control policy. The complexity of the resulting reward landscape leads simple gradient-based methods to become trapped in local optima quickly. Instead of directly solving a long-horizon policy training problem, following [36, 30], where long-horizon simulation tasks are truncated into short-horizon simulation, we utilize short-horizon policy learning to achieve stable gradient backpropagation over a short horizon.
III-F Alignment with A Non-Differentiable Simulator
@@
Due to the simplification of our single rigid-body dynamics, the robot state can diverge from its actual one, eventually leading to unrealistic states due to compounding errors. We proposed to align the body state in our differentiable simulation using information from other simulators that use accurate whole-body dynamics. Specifically, we align the robot state in our differentiable simulation with the state information from IsaacGym [21].
We use the following equation to align the robot state.
$\mathbf{x}^{\text{diff}}_{t+1}=\mathbf{x}^{\text{non-diff}}_{t+1}+\alpha*(\mathbf{x}^{\text{diff}}_{t+1}-\mathbf{x}^{\text{diff, detach}}_{t+1})$ (5)
@@
Here, $\mathbf{x}^{\text{diff, detach}}_{t+1}$ and $\mathbf{x}^{\text{diff}}_{t+1}$ represent the same robot state of our differentiable simulator at time step $t+1$; hence, they share the same value. The word detach indicates that $\mathbf{x}^{\text{diff, detach}}_{t+1}$ is detached from the computational graph for automatic differentiation. Therefore, we can reset the robot state using the value from the non-differentiable simulation during forward simulation: $\mathbf{x}^{\text{diff}}_{t+1}=\mathbf{x}^{\text{non-diff}}_{t+1}+\alpha*\bm{0}$. During backpropagation, the gradient of the state at a given time $t$ is computed as follows
$\partial\mathbf{x}^{\text{diff}}_{t+1}/\partial\mathbf{x}^{\text{diff}}_{t}=\bm{0}+\alpha*\partial\mathbf{x}^{\text{diff}}_{t+1}/\partial\mathbf{x}^{\text{diff}}_{t}.$ (6)
Here, $0<\alpha\leq 1$ is used to decay the gradient.
@@
The non-differentiable simulator can simulate complex contact dynamics and is used to align our simplified model, thereby ensuring that our differentiable training pipeline remains grounded in realistic dynamics. At the same time, our approach benefits from the simplified differentiable simulator, which offers smooth gradients for backpropagation. Fig. 2 shows the computational graph of the forward propagation and backpropagation using alignment.
III-G Differentiable Loss Function
A fundamental difference between policy training using differentiable simulation and reinforcement learning is that the loss function has to be differentiable when using differentiable simulation. On the one hand, RL allows the direct optimization of non-differentiable rewards, such as binary rewards $0/1$. On the other hand, differentiable simulation requires a smooth differentiable cost function to provide learning signals for the desired control inputs.
@@ III-G Differentiable Loss Function
We formulate a loss function tailored for velocity tracking, where the main objective is to follow a specified velocity command, denoted as $\mathbf{v}^{\text{ref}}$. Additionally, we maintain the robot’s body height, represented by $p_{z}$. To enhance the robustness of this system, we incorporate several regularization terms: one to mitigate large angular velocity, thus controlling the body rate; another to limit the output action, preventing large control actions; and a term aimed at stabilizing the robot’s orientation using projected gravity vector.
$l(\mathbf{x}_{t},\mathbf{u}_{t})$ $=$ $b\|\mathbf{v}-\mathbf{v}^{\text{ref}}\|^{2}+a\|p_{z}-p_{z}^{\text{ref}}\|+c\|\bm{\omega}\|^{2}$ (7)
$+d\|\mathbf{u}\|^{2}+f\|\mathbf{g}_{\text{proj}}\|+g\|\mathbf{p}_{\text{foot}}-\mathbf{p}_{\text{foot}}^{\text{ref}}\|^{2}$
Finally, we use a foot position loss function to provide a learning signal for the swing legs. This loss is critical for the swing leg since it contains information about the motor position. The swing trajectory is computed by fitting a quadratic polynomial over the lift-off, mid-air, and landing position of each foot, where the lift-off position is the foot location at the beginning of the swing phase, the landing position is calculated using the Raibert Heuristics [28].
IV Experimental Setup
@@
TABLE I: Policy observation and action.
[…]
V Experimental Results
V-A Learning to Walk with One Robot
This section explores whether a robot can learn to walk in a single simulation environment without parallelization. We design a simple velocity tracking task where the robot is required to follow a constant velocity in the $x$-axis, e.g., $v_{x}=0.2~{}m/s$. At each training iteration, we simulate 24 times steps with one single robot. Hence, each training iteration contains only 24 data points. We trained the policy for a total 1000 iterations, which corresponded to 24,000 data points and 4 minutes of real-time experiences in total. The result is given by the learning curve in Fig. 3. Surprisingly, despite extremely limited data, our policy successfully learns to walk after minutes of training. As a comparison, model-free reinforcement learning failed to yield positive outcomes under the same condition.
V-B Learning Diverse Walking Skills on Challenging Terrains
[…]
Fig. 5 shows the learning curves using different numbers of robots for policy training. We compare the learning performance of our method with a model-free RL algorithm (PPO) [31]. Given minimal samples, e.g., only four robots, RL has a plodding convergence speed and does not learn meaningful walking skills, e.g., the policy constantly falls after a few simulation steps. In contrast, differentiable simulation achieves much higher rewards and can acquire useful walking skills, albeit with relatively low success rates.
As the number of robots increases, the performance for both algorithms improved. Notably, the performance improvement for RL is much more significant than our approach. This indicates the that the zero-order gradient estimates used by reinforcement learning is generally inaccurate and requires much more samples to achieve stable training. On the contrary, the first-order gradient estimates used by our differentiable simulation can have very stable and accurate gradients even given very limited simulation samples. Fig. 5 demonstrates diverse walking skills over challenging terrains using a blind policy trained via differentiable simulation.
Figure 4: Learning to walk on challenging terrains, reinforcement learning versus differentiable simulation. Differentiable simulation achieves faster and more stable convergence, given limited simulation samples.
Additionally, we compare the training wall-clock time and the performance of the resulting policies between the two methodologies. Differentiable simulation often involves the creation of a complex computational graph to facilitate gradient computation, especially for backpropagation through time. As the simulation horizon extends, the length of the computational graph can increase proportionally, leading to a substantial increase in the total training time. This increase can diminish its benefits in situations where gathering a large number of samples is both cheap and fast.
Robots Final Reward Training Time [min]
PPO Ours PPO Ours
4 $-25.82\pm 2.16$ $\mathbf{-3.83\pm 3.01}$ $\mathbf{7.27\pm 0.12}$ $7.28\pm 0.12$
16 $-8.07\pm 1.30$ $\mathbf{-2.03\pm 0.69}$ $7.54\pm 0.06$ $\mathbf{7.50\pm 0.08}$
64 $-6.42\pm 0.29$ $\mathbf{-2.49\pm 0.37}$ $8.46\pm 0.16$ $\mathbf{8.23\pm 0.07}$
1024 $-2.80\pm 0.13$ $\mathbf{-2.69\pm 0.06}$ $\mathbf{11.44\pm 0.13}$ $11.55\pm 0.09$
TABLE II: Learning to walk on challenging terrains, reinforcement learning versus differentiable control. We compare the final reward and the total training walk-clock time.
Figure 5: Diverse locomotion skills over challenging terrains using a blind policy trained via differentiable simulation.
As presented in Table II, a differentiable simulation achieves comparable overall training time. This efficiency is achieved by employing a truncated simulation timestep, which results in a smaller computational graph and enables fast and effective backpropagation. RL requires marginally longer training time due to its mini-batch and multi-epoch training scheme, in contrast, differentiable simulation only uses one gradient update. Differentiable simulation achieved overall higher reward during evaluation.
V-C On the Importance of Non-differentiable Terminal Penalty

### v3

@@ 2.3 Backpropagation Through Time
$\nabla_{\theta}\mathcal{L}_{\theta}$ $=\frac{1}{N}\sum_{k=0}^{N-1}\left(\sum_{i=1}^{k}\frac{\partial l_{k}}{\partial x_{k}}\prod_{j=i}^{k}\underbrace{\left(\frac{\partial x_{j}}{\partial x_{j-1}}\right)}_{\text{differentiable dynamics}}\frac{\partial x_{i}}{\partial\theta}+\frac{\partial l_{k}}{\partial u_{k}}\frac{\partial u_{k}}{\partial\theta}\right),$ (4)
@@
where the matrix of partial derivatives $\partial x_{j}/\partial x_{j-1}$ is the Jacobian of the dynamical system $f$. Therefore, we can compute the policy gradient directly by backpropagating through the differentiable physics model and a loss function $l_{k}$ that is differentiable with respect to the system state and control inputs. A graphical model for gradient backpropagation in policy learning using differentiable simulation is given in Figure 1. Due to the usage of multiplication $\prod$, there are two potential issues in using Eq. (4) for policy gradient: 1) gradient vanishing or exploding, 2) long computation time. We tackle these two problems via short-horizon policy training.
2.4 State Alignment with A Non-Differentiable Simulator
@@
Due to the simplification of our single rigid-body dynamics and the decoupling of the simulation space, the robot state can diverge from its actual one. Over time, even minor discrepancies can accumulate, leading to unrealistic states and ultimately causing the failure of policy training. We proposed to align the body state in our differentiable simulation using information from other simulators that use accurate whole-body dynamics. Specifically, we align the robot state in our differentiable simulation with the state information from IsaacGym [19].
@@
We use the following equation to align the robot state $\mathbf{\hat{x}}^{\text{diff}}_{t+1}=\mathbf{x}^{\text{non-diff}}_{t+1}+\alpha*(\mathbf{x}^{\text{diff}}_{t+1}-\mathbf{x}^{\text{diff, detach}}_{t+1}).$ Here, $\mathbf{x}^{\text{diff, detach}}_{t+1}$ and $\mathbf{x}^{\text{diff}}_{t+1}$ represent the same robot state of our differentiable simulator at time step $t+1$; hence, they share the same value. The word detach indicates that $\mathbf{\hat{x}}^{\text{diff, detach}}_{t+1}$ is detached from the computational graph for automatic differentiation. Therefore, we can reset the robot state using the value from the non-differentiable simulation during forward simulation: $\mathbf{\hat{x}}^{\text{diff}}_{t+1}=\mathbf{x}^{\text{non-diff}}_{t+1}+\alpha*\bm{0}$. During backpropagation, the gradient of the state at a given time $t$ is computed as follows $\partial\mathbf{\hat{x}}^{\text{diff}}_{t+1}/\partial\mathbf{x}^{\text{diff}}_{t}=\bm{0}+\alpha*\partial\mathbf{x}^{\text{diff}}_{t+1}/\partial\mathbf{x}^{\text{diff}}_{t}.$ Here, $0<\alpha\leq 1$ is used to decay the gradient.
@@
The non-differentiable simulator can simulate complex contact dynamics and is used to align our simplified model, thereby ensuring that our differentiable training pipeline remains grounded in realistic dynamics. At the same time, our approach benefits from the simplified differentiable simulator, which offers smooth gradients for backpropagation. Figure 2 shows the computational graph of the forward propagation and backpropagation using state alignment.
2.5 Short-Horizon Policy Training
Although smoothed physical models (e.g., single rigid-body dynamics) improve the local optimization landscape, the complexity of the optimization problem escalates significantly in long-horizon problems involving extensive concatenation of simulation steps. The situation further deteriorates when the actions within each step are interconnected through a nonlinear and nonconvex neural network control policy. The complexity of the resulting optimization landscape can make gradient-based methods to become trapped in local optima quickly. Instead of directly solving a long-horizon policy training problem, following [13, 1], where long-horizon simulation tasks are truncated into short-horizon simulation, we utilize short-horizon policy learning to achieve stable gradient backpropagation over a short horizon. Mathematically, short-horizon policy learning involves truncating a long-horizon trajectory of length $N$ into several shorter segments to compute the policy gradient shown in Eq. (4). For instance, for a trajectory of length 240, we divide it into ten segments of $N=24$. However, it’s important to note that while a shorter horizon could simplify computation and facilitate optimization, it may also limit the model’s ability to anticipate future events, introducing bias into the gradient calculation. This bias could, in turn, affect the types of tasks the model can effectively learn.
2.6 Differentiable Loss Function
@@
A fundamental difference between policy training using differentiable simulation and reinforcement learning is that the loss function has to be differentiable when using differentiable simulation. RL allows the direct optimization of non-differentiable rewards, such as binary rewards $0/1$. On the contrary, differentiable simulation requires a smooth differentiable cost function to provide learning signals for the desired control inputs. We formulate a differentiable loss function $l(\mathbf{x}_{t},\mathbf{u}_{t})$ tailored for velocity tracking, where the main objective is to follow a specified velocity command, denoted as $\mathbf{v}^{\text{ref}}$. Additionally, we maintain the robot’s body height, represented by $p_{z}$. To enhance the robustness of this system, we incorporate several regularization terms: one to mitigate a large angular velocity, thus controlling the body rate $\bm{\omega}$; another to limit the output action $\mathbf{u}$, preventing large control actions; and a term aimed at stabilizing the robot’s orientation using projected gravity vector $\mathbf{g}_{\text{proj}}$. The loss function is defined as
$l(\mathbf{x}_{t},\mathbf{u}_{t})$ $=$ $a_{1}\|\mathbf{v}-\mathbf{v}^{\text{ref}}\|^{2}+a_{2}\|p_{z}-p_{z}^{\text{ref}}\|+a_{3}\|\bm{\omega}\|^{2}$ (5)
$+a_{4}\|\mathbf{u}\|^{2}+a_{5}\|\mathbf{g}_{\text{proj}}\|+a_{6}\|\mathbf{p}_{\text{foot}}-\mathbf{p}_{\text{foot}}^{\text{ref}}\|^{2},$
where $\mathbf{p}_{\text{foot}}$ is the foot position.
3 Experimental Results
3.1 A Toy Example: Control of A Double Integrator
Figure 3: Control of a double integrator using optimal control, reinforcement learning, and differentiable simulation. (left): Learning curves. (middle): Trajectories of different control policies. We initialize the system at the same states for all methods. (right): DiffSim achieves control commands close to optimal control.
Inspired by [20], where the author studied the connection between reinforcement learning and optimal control using a toy double integrator example, we begin by examining the same toy example: control of a double integrator, which is a fundamental problem in system and control theory. The double integrator is a second-order control system that models simple point mass dynamics in one-dimensional space. The state variables are position $x$ and velocity $\dot{x}$, with control inputs $u=\ddot{x}$. The objective is to drive the system state from any initial point to the origin, $[x,\dot{x}]=[0,0]$. Therefore, this problem can be solved by the Linear quadratic regulator (LQR), which can provide an optimal solution. We train a 2-layer multiplayer perception using backpropagation through time via differentiable simulation and the proximal policy optimization (PPO) method [10]. We show that the policy trained using differentiable simulation (DiffSim) achieves nearly optimal performance within a handful of training iterations and samples. In contrast, even when scaling the number of simulation environments to 1024 and training the policy with considerably more iterations, PPO fails to achieve the same level of control performance as LQR or DiffSim. This toy example suggests that scaling might be insufficient for PPO to achieve an optimal solution.
3.2 Learning to Walk with One Robot
This section explores whether a robot can learn to walk without parallelization and with very limited data points at each training iteration. We design a simple velocity-tracking task where the robot is required to follow a constant velocity in the $x$-axis, e.g., $v_{x}=0.2~{}m/s$. At each training iteration, we simulate 24-time steps with one single robot. Hence, each training iteration contains 24 data points. We trained the policy for a total of 1000 iterations. The result is given by the learning curve in Figure 4. Despite very limited data, our policy successfully learns to walk after minutes of training. As a comparison, PPO failed to achieve useful locomotion skills under the same condition. By providing first-order gradients with low variance, differentiable simulations offer a fundamental advantage: they allow for more efficient and reliable optimization of the control policy. This is because the gradient provides a clear direction for updating parameters, and the low variance ensures that this direction is consistent and reliable, even when data is limited.
3.3 Learning Diverse Walking Skills on Challenging Terrains
Figure 5: Learning to walk on challenging terrains, reinforcement learning versus differentiable simulation. Differentiable simulation exhibits significant advantages over PPO in terms of sample efficiency and learning stability. After training in simulation, the policy can be transferred to the real world without fine-tuning.
[…]
As the number of robots increases, the performance of both algorithms improves. Notably, the performance improvement for RL is much more significant than our approach. This indicates that the zeroth-order gradient estimates used by reinforcement learning are generally inaccurate and require many more samples to achieve stable training. On the contrary, the first-order gradient estimates used by our differentiable simulation can have very stable and accurate gradients, even given very limited simulation samples. Figure 5 demonstrates diverse walking skills over challenging terrains using a blind policy trained via differentiable simulation.
[…]
Robots Final Reward Total Training Time [min]
PPO Ours PPO Ours
4 $-25.82\pm 2.16$ $\mathbf{-3.83\pm 3.01}$ $7.27\pm 0.12$ $7.28\pm 0.12$
16 $-8.07\pm 1.30$ $\mathbf{-2.03\pm 0.69}$ $7.54\pm 0.06$ $7.50\pm 0.08$
64 $-6.42\pm 0.29$ $\mathbf{-2.49\pm 0.37}$ $8.46\pm 0.16$ $8.23\pm 0.07$
1024 $-2.80\pm 0.13$ $\mathbf{-2.69\pm 0.06}$ $11.44\pm 0.13$ $11.55\pm 0.09$
Table 1: A comparison of final reward and total training time. We run both methods for the same number of training iterations and collect the same amount of samples.
Additionally, we compare the total training wall-clock time and the final performance of the resulting policies. Differentiable simulation often involves the creation of a complex computational graph to facilitate gradient computation. As the simulation horizon extends, the length of the computational graph can increase proportionally, leading to a substantial increase in the total training time. This increase can diminish its benefits in situations where gathering a large number of samples is both cheap and fast. Therefore, we ask the question: Can differentiable simulation be effectively applied to large-scale settings, for example, thousands of environments? To ensure a fair comparison, we ran both PPO and our training framework for the same number of training iterations, namely 1000 iterations. We compare the total training time, which includes the time for simulating dynamics and updating policies. As shown in Table 1, our results indicate that differentiable simulation can be as time efficient as PPO in large-scale environments despite the requirement of backpropagation through time (BPTT)—a process known for its computational demands. However, notice that the training time to reach a certain reward is much lower with our method and requires less robot and data. For example, our method requires only roughly 64 robots and 200 training iterations to solve this task, while PPO still performs less after 1000 training iterations with 1024 robots.
4 Related Work
Previous efforts in leveraging differentiable simulation to robot control can be broadly categorized into two classes: improving the simulator itself and enhancing the associated optimization algorithms. The first kind has led to the development of several general-purpose differentiable simulators, including Brax [11], Nimble [22], NeuralSim [23], and Dojo [24]. These simulators enable the integration of robot models directly into policy optimization, allowing neural network control policies to be trained using more mathematically precise analytical gradients. However, as noted in [18, 25], in scenarios where the underlying system exhibits chaotic behavior—where small perturbations in initial conditions lead to significantly divergent states, such as in robotic simulations involving contact dynamics—these gradients can become unmanageable, leading to divergence in optimization processes. Our work highlights the benefits of employing well-behaved proxy dynamics as an alternative to the true robot dynamics involving contact. Such a design choice has its roots in model predictive control, which often leverages a simplified model for the optimization to control complex robotic systems [16, 17, 26].
Besides developing a better simulator, various methods have been proposed to improve policy training algorithms. For example, [25] proposes an $\alpha$-order gradient estimator that aims to combine the efficiency of first-order estimates with the robustness of zeroth-order methods. SHAC [13] addresses local minima issues by learning a smooth value function and mitigates vanishing/exploding gradients through truncated backpropagation. Gradient norm clipping strategies have also proven effective in managing the exploding gradients issue [27]. Furthermore, since differentiating through chaotic dynamics is challenging, it is often useful to learn an approximation of these functions and use the approximation for gradient computation. This approach commonly has a connection with model-based reinforcement learning [28, 29, 30, 31, 32], which involves learning a predictive model of the environment, followed by training a controller. Similar to [13], our method leverages a truncated backpropagation to facilitate optimization. Our method could be improved further by combining other ideas, such as learning a critic function or clipping the gradients.
5 Limitations
A key advantage of model-free RL lies in its ability to directly optimize task-level rewards [33], especially non-differentiable rewards. In contrast, differential simulation depends on the backpropagation of information from its objective function, making it challenging to explore novel solutions guided by task objectives. Consequently, our system requires specifying foot positions using the Raibert heuristic [34], as it cannot explore foot placement motion through velocity tracking loss alone. In addition, RL’s flexibility allows for more robust performance enhancements, such as implementing simple termination penalties to discourage falling behaviors or employing constant survival rewards to encourage continued operation. In addition, differentiable simulations are often tailored to specific optimization tasks and might not generalize well across different problems. It requires specific domain knowledge in both building the simulation pipeline and formulating the problem. To generalize our approach to other tasks, additional engineering efforts are required.
6 Conclusion
This work proposes a new perspective on learning quadruped locomotion policies using differentiable simulation by integrating a simplified rigid-body model with a non-differentiable model for whole-body dynamics. Our key insight is that a neural network control policy can be optimized directly by backpropagating the differentiable loss function through a simplified model while simultaneously aligning the robot’s state using a more accurate whole-body dynamics model. Future work should address several limitations of the proposed framework, such as removing the assumption of gait scheduling and the Raibert heuristic for foot contact sequence and foot position planning. Additionally, more sim-to-real analysis is helpful in evaluating the benefits of using a simplified model compared to the true but chaotic gradients or to the policy trained via domain randomization. Finally, we should demonstrate the true advantages of stable training introduced by the first-order gradient by tackling more challenging tasks, such as large-scale training and vision-based control.
7 Acknowledgment
This work was supported by the European Union’s Horizon Europe Research and Innovation Programme under grant agreement No. 101120732 (AUTOASSESS) and the European Research Council (ERC) under grant agreement No. 864042 (AGILEFLIGHT). The authors especially thank Yuang Zhang for his help with the differentiable simulation development and for the useful discussion, Seungwoo Hong for the useful discussion about rigid-body dynamics and nonlinear model predictive control, and Nico Messikommer for his valuable feedback. Also, the authors thank Se Hwan Jeon, Ho Jae Lee, Matthew Chignoli, Steve Heim, Elijah Stanger-Jones, and Charles Khazoom for their helpful support.
Appendix
Appendix 1 Double Integrator
A double integrator system is characterized by its position and velocity. The control input $u$ directly controls the acceleration of the system. For a fair comparison, we formulate the problem as a discrete-time finite-horizon optimal control problem. The discrete-time state space representation of a double integrator is
$x_{k+1}=Ax_{k}+Bu_{k}$
where $A=\begin{bmatrix}1&d_{t}\\ 0&1\end{bmatrix}$ and $B=\begin{bmatrix}d_{t}^{2}/2\\ dt\end{bmatrix}$. Here, $d_{t}$ is the simulation time step. The objective is to minimize a quadratic loss function
$J$ $=x_{N}^{T}Q_{f}x_{N}+\sum_{k=0}^{N-1}(x_{k}^{T}Qx_{k}+u_{k}^{T}Ru_{k})$
$Q$ $=Q_{f}=\begin{bmatrix}1&0\\ 0&1\end{bmatrix},R=[1].$
This is a discrete-time finite-horizon Linear–Quadratic Regulator problem, where the optimal control law can be found using dynamic programming
$u_{k}^{\ast}$ $=-(R+B^{T}P_{k+1}B)^{-1}B^{T}P_{k+1}Ax_{k},\quad k=0,1,\cdots,N-1,$
where $P_{N}=Q_{f}$ and $P_{k}$ can be found from the Riccati recursion.
Appendix 2 Experimental Setup
@@
Table 2: Policy observation and action for quadruped locomotion.
Appendix 3 Foot Trajectory Planning
We use the foot position loss $\|\mathbf{p}_{\text{foot}}-\mathbf{p}_{\text{foot}}^{\text{ref}}\|^{2}$ to provide a learning signal for the swing legs. This loss term is critical for the swing leg since it contains information about the motor position. Following prior works [35, 36], the swing leg trajectory can be computed by fitting a quadratic polynomial over the lift-off $\mathbf{p}^{\text{lift}}_{\text{foot}}$, mid-air $\mathbf{p}^{\text{air}}_{\text{foot}}$, and landing position $\mathbf{p}^{\text{land}}_{\text{foot}}$ of each foot, where the lift-off position is the foot location at the beginning of the swing phase, the landing position $\mathbf{p}^{\text{land}}_{\text{foot}}$ is calculated using the Raibert Heuristics [34], which is expressed as the following function
$\mathbf{p}^{\text{land}}_{\text{foot}}=\mathbf{p}^{\text{hip}}_{\text{foot}}+\mathbf{v}^{\text{CoM}}T_{\text{stance}}/2$
where $T_{\text{stance}}$ is the expected time the foot will spend on the ground, $\mathbf{p}^{\text{hip}}_{\text{foot}}$ is the location on the ground beneath the robot’s hip, $\mathbf{v}^{\text{CoM}}$ is the body velocity projected on the $xy$-plane. The desired contact state of each leg. e.g., swing or stance, is determined via a gait generator. The gait is modulated by a phase variable $\phi\in[0,2\pi]$. The phase is defined through a dynamic function $\phi_{t+1}=\phi_{t}+2\pi f\Delta t$, where $f$ is the stepping frequency. We can design different locomotion patterns by adapting the stepping frequency $f$ and the phase difference between each leg.
Appendix 4 On the Importance of Non-differentiable Terminal Penalty